OpenStudy (anonymous):

What is the x-intercept of the line given by the equation 2x+3y=12

OpenStudy (anonymous):

3y=-2y+12 \[y=-\frac{ 2 }{ 3 }x+3\]

OpenStudy (anonymous):

whats the x-intercept?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

no remember y=mx+b y=-2/3x+3 x= x-intercept b=y-intercept

OpenStudy (tkhunny):

2x+3y=12 Divide by 12 \(\dfrac{x}{6} + \dfrac{y}{4} = 1\) Rather staring at you in this, The Intercept Form!

OpenStudy (anonymous):

Im so lost with this

OpenStudy (anonymous):

when finding the intercepts you need to put it y=mx+b for y=-2/3x+4 what number is with x?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

is the 2 alone?

OpenStudy (anonymous):

-2

OpenStudy (tkhunny):

When finding the y-intercept, do that. When finding both intercepts, simultaneously, use the Intercept Form. Substitute y = 0 for the x-intercept. Substitute x = 0 for the y-intercept.

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 3 }x+4\]

OpenStudy (anonymous):

im doing it the wrong way my mistake im finding slope tkhunny is right

OpenStudy (anonymous):

2x+3(0)=12 2x=12 whats 12 divided by 2

OpenStudy (anonymous):

6

OpenStudy (anonymous):

x=6

OpenStudy (tkhunny):

Remember this? \(\dfrac{x}{6}+\dfrac{y}{4}=1\) Check out that "6" under the "x".

OpenStudy (anonymous):

looks pretty good

OpenStudy (zzr0ck3r):

this would have been just as easy to leave it in the form it was, let y = 0, and solve for x

OpenStudy (zzr0ck3r):

i.e. solve 2x=12

OpenStudy (anonymous):

What is the y-intercept of the line given by the equation y=2x-9

OpenStudy (anonymous):

2(0)-9 solve for that

OpenStudy (zzr0ck3r):

for the equation \(y=mx+b\) \((0,b)\) is the y-intercept

OpenStudy (anonymous):

so y always equals 0?

OpenStudy (zzr0ck3r):

x = 0

OpenStudy (zzr0ck3r):

\((0,a)\) is the form of the y intercept \((a,0)\) is the form of the x intercept

OpenStudy (anonymous):

if you're looking for x interecept plug 0 in for y if you're looking for y intercept plug 0 in for x

OpenStudy (zzr0ck3r):

the intercept is not a number though, its the ordered pair

OpenStudy (anonymous):

y=18

OpenStudy (zzr0ck3r):

?